OpenStudy (anonymous):

(I checked this is the last one my nephew thought he had two more) What is the solution to the equation? 3x – 6 = –2x + 9 A. 5 over 2 B. 3 over 5 C. 3 D. 15 (If this is an easy one don't judge me I own a restaurant I never had to use this!!)

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

You want to solve for x so:

OpenStudy (anonymous):

you're solving the variables on both sides, therefore you have to take away the variables first.

OpenStudy (anonymous):

Wouln't it say solve for x if I needed to solve for x?

OpenStudy (anonymous):

3x - 6 = -2x + 9 Get X over on one side Add 2x from both sides 5x - 6 = 9 5x = 15 x = 15/5 x = 3

OpenStudy (anonymous):

its the only variable there @BadMath021402

OpenStudy (anonymous):

If there's x in it, then you know that you will be wanting to solve for x. It's the only possibility. :)

OpenStudy (anonymous):

Thx. I'll fan both of you

OpenStudy (anonymous):

Thanks! :) Hope it helped!